OpenStudy (anonymous):

use the power and root laws to simplify and then evaluate. \[\log_{3}8^{25} \]

OpenStudy (lollylau):

messy.

OpenStudy (anonymous):

oops sorry guys! i miswrote it!

OpenStudy (lollylau):

"confused"

OpenStudy (anonymous):

\[\log_{3}\sqrt[3]{9} \]

OpenStudy (anonymous):

\[\sqrt[m]n=n^\frac{1}{m}\] \[\log_aa^b=b\]

OpenStudy (anonymous):

use this rules

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what about in these cases? \[\log_{5} 5\sqrt{5}\] and \[\log_{8} 2^{3 \over 2}\]